OpenStudy (anonymous):

i need help compute discrimnant equation for 10^2+x-3=0 can some one show

OpenStudy (anonymous):

can you help

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

\[10x^2+x-3=0\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

for an equation of the form A*x2 + B*x + C = 0 the discriminant is given as D = B^2 - 4*A*C

OpenStudy (anonymous):

put the value of A B and C

OpenStudy (anonymous):

is it ok @shans12

OpenStudy (anonymous):

so it would be x^2-4*10*3

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so what would x^2 be 1

OpenStudy (anonymous):

you take x=1

OpenStudy (anonymous):

im lost could you please show me step by step

OpenStudy (anonymous):

just wait

OpenStudy (anonymous):

The discriminant is a number that can be calculated from any quadratic equation

OpenStudy (anonymous):

A quadratic equation is an equation that can be written as ax ² + bx + c where a ≠ 0

OpenStudy (anonymous):

The discriminant in a quadratic equation is found by the following formula and the discriminant provides critical information regarding the nature of the roots/solutions of any quadratic equation. discriminant= b² − 4ac

OpenStudy (anonymous):

r u getting @shans12

OpenStudy (anonymous):

LOL NO IM NOT I FEEL STUPID

OpenStudy (anonymous):

then now u plz tell the basic formula so that i could xceed

OpenStudy (mayankdevnani):